OpenStudy (anonymous):

3 7/8 divided by 3 2/9 write as a mixed number I don't know how to set up?

OpenStudy (anonymous):

(37/8)*(32/9)=333/256

OpenStudy (anonymous):

then u divide them and turn into a mixed number?

OpenStudy (anonymous):

77 1---- answer matching? 256

OpenStudy (anonymous):

ya this will be the answer.. in mixed number

OpenStudy (anonymous):

medel :)

OpenStudy (anonymous):

eliseshamberger ??? matching??

OpenStudy (anonymous):

sorry I was figuring it hot ty for showing me how to set up the problem

OpenStudy (anonymous):

out not hot lol

OpenStudy (anonymous):

\[\frac{3+\frac{7}{8}}{3+\frac{2}{9}}=\frac{\frac{31}{8}}{\frac{29}{9}}=\frac{31}{8}\frac{9}{29}=\frac{279}{232} \]

OpenStudy (anonymous):

\[\frac{279}{232}=1+\frac{\text{Mod}[279,232]}{232}=1+\frac{47}{232}=1\frac{47}{232}\]